# Counterfactual Fairness through Transforming Data Orthogonal to Bias

### v1

@@ 3.2. Achieving counterfactual fairness via data decorrelation
We first demonstrate that $\hat{Y}$ is counterfactually fair, which is achieved when the model’s predictions for $Y$ are not influenced by the sensitive variable $B$. In the following proof, we focus on the case of a binary predictor $\hat{Y}$ for simplicity, noting that our findings can be seamlessly applied to predictors that yield continuous outcomes. This simplification allows us to establish fairness by showing that the expected outcomes are equivalent, which for a Bernoulli random variable, also indicates a distributional equivalence. Following the well-established “Abduction-Action-Prediction” method from (Pearl, 2009), the conditional expectation of $\hat{Y}_{\mathbf{b}^{\prime}}$ given $B=\mathbf{b}^{*},A=\mathbf{a}^{*}$ can be written as:
(3) $~{}\mathbb{E}(\hat{Y}_{\mathbf{b}^{\prime}}|B=\mathbf{b}^{*},A=\mathbf{a}^{*})$
@@
Next, we prove that $A$ is independent of $B$ when they are uncorrelated under Assumption 3.2, which is a commonly accepted statistical result. Consider the mean vectors for $A$ and $B$ as $\boldsymbol{\mu}_{A}$ and $\boldsymbol{\mu}_{B}$, respectively, with covariance matrices $\Sigma_{A}$, $\Sigma_{B}$. Recall that when $A$ and $B$ uncorrelated, the covariance matrix $\Sigma$ of $A$ and $B$ is
$\Sigma=\left[\begin{array}[]{cc}\Sigma_{A}&0\\ 0&\Sigma_{B}\end{array}\right]\text{~{}and~{}}\Sigma^{-1}=\left[\begin{array}[]{cc}\Sigma_{A}^{-1}&0\\ 0&\Sigma_{B}^{-1}\end{array}\right].$
@@
Substituting $\Sigma$ above into the joint probability density function of $A$ and $B$, we have
(4) $\mathbb{P}(\mathbf{a},\mathbf{b})=$ $~{}\frac{1}{\sqrt{(2\pi)^{q+p}|\Sigma|}}\cdot$
@@
$~{}\exp\left(-\frac{1}{2}\left[\begin{array}[]{l}\mathbf{a}-\boldsymbol{\mu}_{A}\\ \mathbf{b}-\boldsymbol{\mu}_{B}\end{array}\right]^{\top}\Sigma^{-1}\left[\begin{array}[]{l}\mathbf{a}-\boldsymbol{\mu}_{A}\\ \mathbf{b}-\boldsymbol{\mu}_{B}\end{array}\right]\right)$
$=$ $~{}\frac{1}{\sqrt{(2\pi)^{q+p}|\Sigma_{A}||\Sigma_{B}|}}\cdot$
$~{}\exp\left(-\frac{1}{2}(\mathbf{a}-\boldsymbol{\mu}_{A})^{\top}\Sigma_{A}^{-1}(\mathbf{a}-\boldsymbol{\mu}_{A})\right)\cdot$
$~{}\exp\left(-\frac{1}{2}(\mathbf{b}-\boldsymbol{\mu}_{B})^{\top}\Sigma_{B}^{-1}(\mathbf{b}-\boldsymbol{\mu}_{B})\right)$
$=$ $~{}\mathbb{P}_{A}(\mathbf{a})\mathbb{P}_{B}(\mathbf{b}).$
@@
As we can observe that the joint distribution decomposes into the product of their marginal distributions, hence demonstrating their statistical independence when $A$ and $B$ are jointly normal and uncorrelated.
@@
Finally, establishing $A$ and $B$ as jointly normal and uncorrelated leads to the inference that, given the SCM in (1), predictor $\hat{Y}$ is counterfactually fair.
@@
Theorem 3.3 suggests that achieving counterfactual fairness in the predictor $\hat{Y}$ is possible by decorrelating non-sensitive variables $A$ from sensitive variables $B$. This insight motivates us to develop a data pre-processing algorithm aimed at adjusting the observed data with minimal changes to achieve uncorrelation between non-sensitive and sensitive variables.
It is important to emphasize that Assumption 3.2 is applicable across a broad spectrum of applications where standardization of data is required, ensuring that all variables, both sensitive and non-sensitive, are normalized to have a mean of zero and a variance of $1$. Moreover, our investigations have revealed that this assumption is not always critical for the success of our data pre-processing algorithm. Our approach has demonstrated effectiveness even when applied to data sets that do not meet this criterion. In particular, as elaborated in Section 4, our strategy has proven to be promising in experiments that involve categorical sensitive variables $B$, rather than continuous ones.
3.3. Orthogonal to bias
@@
In this section, we develop a data pre-processing algorithm, termed as Orthogonal to Bias (OB). We first standardize both non-sensitive variables $A$ and sensitive variables $B$, to achieve a normal distribution for each. Then the empirical covariance between $A$ and $B$ can be estimated by
@@
$\text{cov}(A,B)=\mathbb{E}[(A-\mathbb{E}[A])(B-\mathbb{E}[B])]\approx\frac{1}{n}\sum_{i=1}^{n}\mathbf{a}_{i}^{\top}\mathbf{b}_{i}=\left\langle\mathbf{A},\mathbf{B}\right\rangle.$
Given that two variables are uncorrelated if their covariance is zero, orthogonality between observed data $\mathbf{A}$ and $\mathbf{B}$ guarantees uncorrelation. Therefore, the OB algorithm aims to adjust the observed non-sensitive data $\mathbf{A}$ in such a way that it is orthogonal to the observed sensitive data $\mathbf{B}$, while ensuring minimal changes to non-sensitive data $\mathbf{A}$.
@@
Specifically, we follow the idea of Orthogonal to Groups introduced by (Aliverti et al., 2021), and define a rank $k$ approximation of $\mathbf{A}$ as $\widetilde{\mathbf{A}}=\mathbf{S}\mathbf{U}^{\top}$, where $\mathbf{U}=[\textbf{u}_{1},\dots,\textbf{u}_{k}]$ is a $q\times k$ orthonormal matrix and $\mathbf{S}=\{s_{ij}\}$ is a $n\times k$ matrix. The goal is to find a transformed $n\times q$ matrix $\widetilde{\mathbf{A}}$ that is orthogonal to $\mathbf{B}$ with minimal change to matrix, as measured by the Frobenius norm $\|X\|_{\mathrm{F}}=\sqrt{\sum_{i}\sum_{j}x_{ij}^{2}}$. Formally, we aim to solve the following constrained optimization problem:
@@
(5) $\arg\min_{\mathbf{S},\mathbf{U}}{\left\lVert\mathbf{A}-\mathbf{S}\mathbf{U}^{\top}\right\rVert}^{2}_{F}$ ,

### v2

@@ 2.2 Achieving counterfactual fairness via data decorrelation
We first demonstrate that $\hat{Y}$ is counterfactually fair when the model’s predictions for $Y$ are not influenced by the sensitive variable $B$. In the following proof, we focus on the case of a binary predictor $\hat{Y}$ for simplicity, noting that our findings can be seamlessly applied to predictors that yield continuous outcomes. This simplification allows us to establish fairness by showing that the expected outcomes are equivalent, which for a Bernoulli random variable, also indicates a distributional equivalence. Following the well-established “Abduction-Action-Prediction” method from [14], the conditional expectation of $\hat{Y}_{\mathbf{b}^{\prime}}$ given $B=\mathbf{b}^{*},A=\mathbf{a}^{*}$ can be written as:
@@
Next, we prove that $A$ is independent of $B$ when they are uncorrelated under Assumption 2.2, which is a commonly accepted statistical result. Consider the mean vectors for $A$ and $B$ as $\bm{\mu}_{A}$ and $\bm{\mu}_{B}$, respectively, with covariance matrices $\Sigma_{A}\in\mathbb{R}^{q\times q}$, $\Sigma_{B}\in\mathbb{R}^{p\times p}$. Recall that when $A$ and $B$ are uncorrelated, the cross-covariance matrix between $A$ and $B$ is $\Sigma_{[A,B]}\in\mathbb{R}^{q\times p}=\mathbf{0}$. Therefore, the covariance matrix $\Sigma_{[A,B]}$ of joint distribution of $[A,B]$ is
$\Sigma_{[A,B]}=\left[\begin{array}[]{cc}\Sigma_{A}&0\\ 0&\Sigma_{B}\end{array}\right]\text{~{}and~{}}\Sigma_{[A,B]}^{-1}=\left[\begin{array}[]{cc}\Sigma_{A}^{-1}&0\\ 0&\Sigma_{B}^{-1}\end{array}\right].$
@@
Substituting $\Sigma_{[A,B]}$ above into the joint probability density function of $A$ and $B$, we have
@@
$\mathbb{P}(\mathbf{a},\mathbf{b})=$ $~{}\frac{1}{\sqrt{(2\pi)^{q+p}|\Sigma_{[A,B]}|}}\cdot\exp\left(-\frac{1}{2}\left[\begin{array}[]{l}\mathbf{a}-\bm{\mu}_{A}\\ \mathbf{b}-\bm{\mu}_{B}\end{array}\right]^{\top}\Sigma_{A,B}^{-1}\left[\begin{array}[]{l}\mathbf{a}-\bm{\mu}_{A}\\ \mathbf{b}-\bm{\mu}_{B}\end{array}\right]\right)$
$=$ $~{}\frac{1}{\sqrt{(2\pi)^{q+p}|\Sigma_{A}||\Sigma_{B}|}}\cdot\exp\left(-\frac{1}{2}(\mathbf{a}-\bm{\mu}_{A})^{\top}\Sigma_{A}^{-1}(\mathbf{a}-\bm{\mu}_{A})\right)\cdot$
$~{}\exp\left(-\frac{1}{2}(\mathbf{b}-\bm{\mu}_{B})^{\top}\Sigma_{B}^{-1}(\mathbf{b}-\bm{\mu}_{B})\right)=\mathbb{P}_{A}(\mathbf{a})\mathbb{P}_{B}(\mathbf{b}).$
@@
As we can observe, the joint distribution decomposes into the product of their marginal distributions, hence demonstrating their statistical independence when $A$ and $B$ are jointly normal and uncorrelated.
@@
Finally, establishing $A$ and $B$ as jointly normal and uncorrelated implies that, given the SCM in (1), predictor $\hat{Y}$ is counterfactually fair. ∎
@@
Theorem 2.3 suggests that achieving counterfactual fairness in the predictor $\hat{Y}$ is possible by decorrelating non-sensitive variables $A$ from sensitive variables $B$. This insight motivates us to develop a data pre-processing algorithm aimed at minimally adjust the observed data to remove the correlation between non-sensitive and sensitive variables while preserving strong predictive performance.
It is important to emphasize that Assumption 2.2 is applicable across a diverse array of applications that necessitate data standardization. For example, the SCM configuration depicted in Figure 2(a), for instance, aligns with Assumption 2.2 when $A$ is linear in terms of $U_{A}$, as shown in Appendix A. While the assumption of normality may not always hold perfectly, standardization serves as a pragmatic and effective pre-processing step that aligns real-world data more closely with the required conditions. By utilizing this widely accepted practice, empirical results indicate that our method remains both effective and robust, capable of delivering reliable performance even when applied to real-world datasets that do not fully meet this criterion. In particular, as elaborated in Section 3, our strategy has proven to be promising in experiments that involve both categorical and continuous sensitive variables $B$.
2.3 Orthogonal to bias
@@
Now we introduce our data pre-processing algorithm, termed as Orthogonal to Bias (OB). We assume both non-sensitive variables $A$ and sensitive variables $B$ have a zero mean for each column. This allows us to estimate the empirical covariance between $A$ and $B$ by calculating their inner product:
@@
$\text{cov}{(A,B)}=\mathbb{E}[(A-\mathbb{E}[A])(B-\mathbb{E}[B])]\approx\left\langle\mathbf{A},\mathbf{B}\right\rangle.$
Since two variables are uncorrelated if their covariance is zero, achieving orthogonality between the observed data $\mathbf{A}$ and $\mathbf{B}$ ensures their uncorrelation.
@@
Therefore, the key idea of OB algorithm is to adjust the observed non-sensitive data $\mathbf{A}$ in such a way that it is orthogonal to the observed sensitive data $\mathbf{B}$, while ensuring minimal changes to non-sensitive data $\mathbf{A}$. Specifically, we follow the idea of Orthogonal to Groups introduced by [23], and define a rank $k$ approximation of $\mathbf{A}$ as $\widetilde{\mathbf{A}}=\mathbf{S}\mathbf{U}^{\top}$, where $\mathbf{U}=[\textbf{u}_{1},\dots,\textbf{u}_{k}]$ is a $q\times k$ orthonormal matrix and $\mathbf{S}=\{s_{ij}\}$ is a $n\times k$ matrix with assosiated scores. The goal is to find a transformed $n\times q$ matrix $\widetilde{\mathbf{A}}$ that is orthogonal to $\mathbf{B}$ with minimal change to matrix, as measured by the Frobenius norm $\|X\|_{\mathrm{F}}=\sqrt{\sum_{i}\sum_{j}x_{ij}^{2}}$. Formally, we aim to solve the following constrained optimization problem:
@@
$\arg\min_{\mathbf{S},\mathbf{U}}{\left\lVert\mathbf{A}-\mathbf{S}\mathbf{U}^{\top}\right\rVert}^{2}_{F},\quad\text{s.t.}~{}\left\langle\mathbf{S}\mathbf{U}^{\top},\mathbf{B}\right\rangle=\mathbf{0},~{}\mathbf{U}\in\mathcal{G}_{q,k},$ (4)
where $\mathcal{G}_{q,k}$ is the Grassman manifold [24]. The last constraint helps prevent degeneracy, such as basis vectors becoming identically zero or encountering solutions with double multiplicity. Here we let $\mathbf{U}$ be orthonormal matrix in practice, i.e., $\mathbf{U}^{\top}\mathbf{U}=\mathbf{I}_{k}$.
In the following, we focus on a univariate $\mathbf{b}\in\mathbb{R}^{n}$ for clarity. We note the procedure can be easily extended for multivariate $B$ [23]. The above constrained optimization (4) can be reformulated in terms of Lagrange multipliers $\lambda_{j}$:
$\arg\min_{\mathbf{S},\mathbf{U}}~{}\frac{1}{n}\sum_{i=1}^{n}\lVert\textbf{a}_{i}-\sum_{j=1}^{k}s_{ij}\textbf{u}_{j}\rVert^{2}+\frac{2}{n}\sum_{j=1}^{k}\lambda_{j}\sum_{i=1}^{n}s_{ij}{b}_{i},$ (5)
where $\lambda_{j}\in\mathbb{R}$ denotes the Lagrange multiplier and $b_{i}$ is the $i$-th item of $\mathbf{b}$. Here, the factor $2/n$ is introduced to simplify the expression for the optimal solutions. It ensures that the first-order condition for (5) with respect to $s_{ij}$ involves a common factor of $2/n$, which can then be canceled out during computations. Let $\mathbf{S}^{*}$ and $\mathbf{U}^{*}$ denote the optimal solutions of $\mathbf{S}$ and $\mathbf{U}$, respectively. The reformulated equation (5) gives a closed-form solution:
$\lambda^{*}_{j}=\frac{\left\langle\mathbf{A}\mathbf{u}^{*\top}_{j},\mathbf{b}\right\rangle}{\langle{\mathbf{b}},\mathbf{b}\rangle},\quad\mathbf{U}^{*}=\left[\textbf{u}^{*}_{1},\dots,\textbf{u}^{*}_{k}\right],\quad\mathbf{S}^{*}=\{s^{*}_{ij}\},$ (6)
where $\mathbf{U}^{*}$ consists of the first $k$ right singular vectors of $\mathbf{A}$ and $s_{ij}^{*}=\mathbf{a}_{i}\mathbf{u}^{*\top}_{j}-\lambda^{*}_{j}{b}_{i}$. The pre-processed non-sensitive data matrix can be obtained by $\widetilde{\mathbf{A}}=\mathbf{S}^{*}\mathbf{U^{*}}^{\top}$. Detailed derivations of the closed-form solution (6) can be found in Appendix B.
It is noteworthy that the reconstruction error of $\widetilde{\mathbf{A}}$ equals to that of the standard Singular Value Decomposition (SVD) if there is zero correlation between the sensitive and non-sensitive variables. It is because the additional reconstruction error of $\widetilde{\mathbf{A}}$ relative to SVD with the same rank is proportional to the collinearity between the subspace spanned by $\mathbf{B}$ and the left singular vectors of the data $\mathbf{A}$ [23]. The comparison of reconstruction error between the OB and SVD can be found in Appendix B.
Sparse orthogonal to bias (SOB)
When the number of features $p$ exceeds the number of observations $n$, estimating a low-dimensional structure from high-dimensional data can become numerically unstable [25]. To address this challenge, we introduce a sparse variant of the OB algorithm, referred to as SOB. The SOB imposes an $\ell_{1}$-norm penalty for $\mathbf{U}$ to encourage sparsity and improve numerical stability, in addition to the orthogonality constraints in (4). We define $h$ as the $\ell_{1}$ constraint on $u$. Details of the formulation of SOB can be found in Appendix C. Following derivation by [23, 26], Algorithm 1 illustrates the key steps to implement the SOB algorithm, where $\eta$ represents the minimum change to terminate the iterative optimization process.
Note that with the additional regularization constraints, the solution favors sparsity while satisfying the orthogonal constraint. Therefore, the SOB also achieves counterfactual fairness under SCM framework with Theorem 2.3.
(a) Counterfactual prediction
(b) Predictive error vs counterfactual fairness
Figure 3: Comparison of our approach (OB/SOB) with the primary baseline methods on the synthetic dataset: (a) Counterfactual predictions for 1,000 randomly generated sample points across the sensitive variable, comparing our approach to baseline methods. Our method’s predictions remain flat across the sensitive variable; (b) Comparison of predictive error and counterfactual fairness between our approach and baseline methods. The dashed line represents an exponential function fitted to the results of Adv. The error bars indicate standard deviation. Our method is under the curve and achieves a low MSE while maintaining relatively low CF compared to Adv.
3 Experiments
In this section, we empirically evaluate the performance of our method using three real-world datasets and two simulated datasets. For the discrete scenario where $Y\in\{0,1\}$ and $B\in\{0,1\}$, we utilize two popular datasets: the Adult income dataset [27], with gender as the sensitive variable (male or female) and whether a person earns over \$50K a year as $Y$, and the COMPAS dataset [28], with race as the sensitive variable (Caucasian or non-Caucasian) and two-year recidivism as $Y$. In addition, we explore a synthetic loan decision dataset using similar setup as [10]. For scenarios where both $Y$ and $A$ are continuous, we use the Crime dataset [29], with the ratio of an ethnic group per population as the sensitive variable and violent crimes per population as $Y$. Additionally, we explore a synthetic scenario that allows for further analysis of the relative performances of the approaches. This synthetic scenario involves a pricing algorithm for a fictional car insurance price adapted from [12], following the causal graph in Figure 2(a). Details of the experimental setup can be found in Appendix D.
Table 1: Performance on Crime and Synthetic Insurance Price Data
Dataset Metrics Baselines Compared Methods Ours
ML FTU Adv ($\lambda=0.1$) Adv ($\lambda=0.2$) OB SOB
Crime [29] MSE 0.4751 (0.0336) 0.4711 (0.0101) 0.4674 (0.0425) 0.4511 (0.0503) 0.4534 (0.0110) 0.4491 (0.0155)
CF-metric 0.2353 (0.0729) 0.2467 (0.1633) 0.0095 (0.0057) 0.0066 (0.0029) 0.1047 (0.0373) 0.1051 (0.0328)
Avg. Training Time (s) 68.9089 (7.4695) 62.9821 (6.5349) 136.9746 (4.4359) 131.8830 (4.4037) 69.2535 (5.1429) 80.7566 (7.4400)
Synthetic MSE 0.0008 (0.0009) 0.0000 (0.0000) 0.0658 (0.0731) 0.1062 (0.2174) 0.0054 (0.0001) 0.0054 (0.0002)
CF-metric 0.1414 (0.0008) 0.1418 (0.0002) 0.1422 (0.0361) 0.1189 (0.0409) 0.1309 (0.0023) 0.1296 (0.0020)
Avg. Training Time (s) 36.4220 (13.9463) 32.9465 (7.0428) 55.2336 (2.6324) 63.5752 (0.6989) 32.8576 (8.6382) 37.5425 (10.8821)
Evaluation metrics
We assess the accuracy of decisions using Mean Squared Error (MSE) for continuous value predictions, and Area Under the Curve (AUC) and Accuracy (ACC) for binary classification problems. In evaluating counterfactual fairness for continuous cases, we utilize the CF-Metric employed by [12], which measures the mean change in predictions for a set of random counterfactual samples for each individual in the test set. These counterfactuals are generated using a variational inference model of the Structural Causal Model (SCM), following the same adversarial training process described by [12]. For classification problems, we adopt CF-Metrics from [10], which measure counterfactual fairness by calculating the average change in predicted scores between two groups. This metric is designed to be zero when decisions are counterfactually fair. For a comprehensive comparison of methods, we also include two additional counterfactual fairness metrics: AA Fairness and Equalized Opportunities (EO) Fairness [9]. Both AA Fairness and EO Fairness are only defined for scenarios with discrete sensitive variables. Definitions of these metrics are provided in Appendix D.
Results with continuous sensitive variables
To our knowledge, the approach by [12] is the only other notable attempt to tackle counterfactual fairness with continuous sensitive variables. In our experiments, we included this method (henceforth referred to as Adv) along with a standard machine learning approach (ML) and the Fairness Through Unawareness (FTU) method as baselines. All models were evaluated using a uniform four-layer neural network for prediction tasks.
Figure 3 illustrates the fairness-accuracy trade-off, a well-explored topic in the literature [20, 21, 22]. Moving along the spectrum, Adv appears with progressively increasing $\lambda$ values, reflecting a shift towards greater emphasis on fairness at the expense of accuracy. Our method is located in the lower-left, demonstrating low MSE and CF metrics, with narrow error bars due to its closed-form solution (Appendix B). Furthermore, Table 1 provides detailed numerical results, including mean and standard deviations from 10 repeated experiments, which consistently demonstrate our method’s superior performance in achieving lower MSE and competitive CF-metrics across two datasets with continuous variables. Moreover, Table 1 also highlights our method’s greater time efficiency compared to Adv. Moreover, Table 1 shows that our method is more time-efficient than Adv.
Table 2: Performance comparison on COMPAS data [28]
Metrics Baselines Compared Methods Ours